OpenStudy (anonymous):

please help!! Which of the following shows the correct steps to find the value of 16 to the power of 1 over 4?

OpenStudy (anonymous):

answer choices: 16 ^1/4=2^4)1/4= (2) 4 1/4= 2 16^1/4= (4^4)1/4= (4) 4 1/4= 4 16^1/4= (2^8)1/4= (8) 4 1/4= 4 16^1/4= (8^2)1/4= (2) 2 1/4=8

OpenStudy (anonymous):

thankyou! i was just a bit confused

OpenStudy (anonymous):

an exponent of 1/4 is a fourth root

OpenStudy (anonymous):

here is a website that might help you understand it better: https://www.mathsisfun.com/algebra/exponent-fractional.html

OpenStudy (anonymous):

It is 2, but try solving more problems after checking out the website I sent you.

OpenStudy (anonymous):

okay i will! thanks!

OpenStudy (anonymous):

no problem, bye